From the 1/14th Daily Journal Summaries for the month of April, 1968.
April 22-30, 1968
22 April 1968
GOLDEN DRAGONS continued operations in assigned AO. An airborne FAC working in direct support of Delta Co located a bunker complex consisting of 20-30 bunkers. The bunkers were searched and then destroyed. Negative findings. Alpha Co had movement around their night logger positions with negative contact.
PLANS: Continue security operations around LZ SWINGER. With Alpha Co in the east, Bravo Co in the west, Delta Co to the south.
23 April 1968
GOLDEN DRAGONS continued ooperations in assigned AO. There was negative contact during the day or night.
PLANS: Alpha Co to sweep north to river, Delta Co to move north to cover high ground southwest of LZ SWINGER. Charlie Co to conduct local patrols around LZ SWINGER. Bravo Co to hold present positions and conduct local patrols.
24 April 1968
D/1/22 has enemy contact. Initial 6 NVA now an undetermined force. Mike 84 requests gun ships and to have a AS on stand by. Gladiator 063 has gone down in the contact area.
Gladiator 063 brought down by 7 or 8 rds of enemy small arms fire. Bird completely destroyed. Pilots and gunners O.K. D/1/2 has 1 US WIA.
D/1/22 has 5 US WIA's. 4 not serious, 1 serious. Still receiving sporadic sniper fire. Most of the enemy have fled to the W. Believed to be a company sized element.
D/1/22's contact considered light. Report 1 more US serious WIA. Enemy continues to flee to the west.
GOLDEN DRAGONS continued operations in assigned AO. Alpha Co while sweeping to the north found a large concealed trail network at grid 857579. Negative enemy sighting. Delta Co 1/22 had contact south of the Bn AO. Enemy fled to the west.
PLANS: Bravo and Charlie Co's to move west and sweep high ground west of LZ SWINGER where indications are that enemy supply point might be located. Delta Co to assume responsibility for firebase security.
25 April 1968
GOLDEN DRAGONS continued operations in assigned AO. Bravo and Charlie Co's found several bunker complexes with bunkers built to withstand 750 lb. bombs. They also found a camoflaged road approximately 15 feet wide going N&S. Indications are that chain saws and building equipment were used to complete the road.
PLANS: B&C Co's are to continue to sweep the road and ajoining high ground. A&D Co's to conduct local patrols.
26 April 1968
GOLDEN DRAGONS continued operations in assigned AO. Charlie Co moving north along the road found an old bivouac area for approx. 200 people. Had not been used recently.
PLANS: Bravo and Charlie Co's to conduct CA to the south in support of 1/22 Inf. Alpha Co to move to LZ SWINGER for security mission.
27 April 1968
C/1/14 has one NVA KIA at loc 838913. Ind was wearing green Unif w/steel pot.
GOLDEN DRAGONS continued operations in assigned AO. Alpha Co found 2 tunnels and some partly burned documents. documents found to be NVA clothing records. Bravo and Charlie Co's conducted CA in support of 1/22 INf. Charlie Co had small contact with NVA. Results 1 NVA KIA. C Company OPCON to 1/22 Inf.
PLANS: Conduct CA with Alpha Co to support ground operations of the 1/22 Inf. Delta Co to remain as security for LZ SWINGER.
28 April 1968
D/1/14 had two WIA at loc 845904. Dustoff completed at 1555 hrs. Extent of wounds unk at this time. 1 apranged angle & 1 cut hand.
D/1/14 recd 5 iincoming mortar rds neg causalties. At 1920 C/1/14 7 WIA &1 KIA. Loc 849897
GOLDEN DRAGONS continued operations in support of 1/22 Inf. Alpha Co conducted CA near Charlie Co 1/14 and began to conduct operations with Charlie. Alpha and Bravo Co's revert OPCON to 1/22. Charlie Co while OPCON had 5 NVA KIA and 1 wpn(AK-47) CIA. Delta also OPCON to 1/22. During CA Delta Co received mortar fire on LZ, results 5 WIA. LZ SWINGER received 7 incoming mortar rounds. AS was called and results one large secondary explosion. 1/18 receive OPCON Bravo Co 1/8 Inf for security of LZ SWINGER.
PLANS: Elements of 1/14 to continue support of operations of 1/22 Inf.
29 April 1968
B/1/14 while OPCON to 1-22 rept rect 60mm, 82mm B-40 rockets and flame throwers. Est fire coming from the East believe to be a Bn or larger. Results 1 US KIA, 5 US WIA. 0510 Total 2 US KIA & 11 WIA, loc 828908.
B/1/14 while OPCON 1-22 it's 10 element is recieving small arms fire from the west. Returned fire results 1 NVA KIA & 1 sks. loc 828908.
B/1/14 while OPCON 1-22 has body count(est) of 150 NVA KIA. loc 828908.
1, B/1/14 is in contact at loc 828908. 3, loc 830910 has 40 NVA KIA's 8 wpns, 1 B-40 Launcher, 2 hvy MG, 3 AK-47s, 2 SKS's, numerous rockets, hand grenades, Ind were dressed in k unif w/new web gear & packs.
B/1/14 identified 2 NVA units as 507th 3/7th Inf.
Gladiator dustoff completed dustoff for B/1/14 6 WIA's wound in last night action. PZ LZ Swinger. Completed at 1300 hrs Also requested dustoff for 4 more of B/1/14 WIA's comp 1308. PZ LZ Swinger.
D/1/14 recd 3 incoming mortar rds neg casualties at this time. Results 13 US WIA's 1 US KIA.
Be prepared to move 2 companies from LZ Swinger to 912958(Hill 1042) at 1000 hr 30 April. The remainder of the 1/14 and 2/9 Art. will move to Polei Kleng sometime the same day. Fire Base Swinger will be completely cleared.
GOLDEN DRAGONS saw Bravo Co while OPCON to 1/22 Inf had ground attack on the defensive position. Results 46 NVA KIA (B.C.) 150 NVA KIA (EST) 8 wpns CIA, 2 heavy MG CIA. numerous rockets and hand grenades CIA. Ind NVA were dressed in tan uniforms with new web equipment and packs. All units returned to 1/14 Inf. Bravo 1/8 reverted OPCON to parent unit. During extraction of Delta Co mortar rounds were received on PZ Results 13 US WIA, 1 US KIA. Alpha Co and Bravo Co returned to original positions upon return to LZ SWINGER. Charlie and Delta Co's to secure LZ SWINGER.
PLANS: 1/14 to move from LZ SWINGER to establish new FSB vic coordinates 899933. New FSB to be named LZ ROBERTS.
30 April 1968
GOLDEN DRAGONS planned move was postponed due to lack of aircraft in the Bde. GOLDEN DRAGONS continued to prepare for CA in the morning.
PLANS: CA to new firebase with three companies. Delta Co and the Command group to move to POLEI KLENG to await the clearing of the new FSB.